08-29-2009
August 29, 2009
ISS On-Orbit Status 08/29/09

All ISS systems continue to function nominally, except those noted previously or below. Saturday --- Very short crew rest day.

STS-128/Discovery (ISS-17A) lifted off spectacularly in darkness last night on time (11:59pm EDT) with all systems performing nominally, for rendezvous with ISS tomorrow evening (8/30, Sunday), set to dock at approximately 9:03pm EDT. We are off to another great mission! [The Orbiter is carrying the seven-member crew of Commander Fredrick W. “Rick” Sturckow, Pilot Kevin A. Ford, and Mission Specialists Patrick G. Forrester, Jose M. Hernandez, John D. “Danny” Olivas, Christer Fuglesang & Nicole Passonno Stott. Stott will replace Timothy Kopra as ISS Flight Engineer 2, who returns on 9/10 (nominal) with STS-128 (the last time for an entirely Shuttle-based crew rotation). STS-128 is the 128th space shuttle flight in history, the 4th Shuttle mission in 2009, the 37th flight for Discovery, and the 30th Shuttle flight to the station. Primary payload for the Discovery is the 15,200-lbs Italian-built MPLM (Multipurpose Logistics Module) Leonardo.]

After their wakeup this morning at the regular 2:00am EDT, the ISS crew today has started another sleep/wake cycle shift to synchronize with 17A arrival and docked activities. This resulted in a short workday, with sleeptime beginning already at 10:00am. Next wakeup for FD2 will be tonight at 6:30pm. [See below for a listing of upcoming sleep/wake times.]

Upon wakeup, crewmembers Barratt, Kopra & Thirsk continued their current support of the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight), logging data from their Actiwatches to the HRF-1 (Human Research Facility 1) laptop as part of a week-long session.

FE-3 Romanenko terminated the recharge of the Kelvin-Video battery pack of the BAR/EXPERT experiment initiated yesterday. [Objective of the Russian KPT-12/EXPERT science payload is to measure environmental parameters (temperature, humidity, air flow rate) and module shell surface temperatures behind SM panels and other areas susceptible to possible micro-destruction (corrosion), before and after insolation (day vs. night). The payload uses a remote infrared thermometer (Kelvin-Video), a thermohygrometer (Iva-6A), a heat-loss thermal anemometer/thermometer (TTM-2) and an ultrasound analyzer (AU) to determine environmental data in specific locations and at specific times. Activities include documentary photography with the NIKON D2X camera and flash.]

To prepare for a potential EVA in support of the HTV (H-II Transfer Vehicle) docking, in the US Airlock FE-1 Barratt configured the EMU (Extravehicular Mobility Unit) #3006 with its SCU (Service & Cooling Umbilical) and initiated the standard one-hour scrubbing process on the spacesuit’s cooling water loops, filtering ionic and particulate matter (via a 3-micron filter), then reconfigured the cooling loops and started the ~2hr biocide filtering. Scrubbing termination, disassembly of the EMU water processing kit and stowing the equipment followed. [Loop scrubbing, incl. iodination of the LCVG (Liquid Cooling & Ventilation Garment) for biocidal maintenance is done to eliminate any biomass and particulate matter that may have accumulated in the loops.]

Mike Barratt also performed a quick fit check on the Water ORU (Orbital Replacement Until) of the OGS (Oxygen Generator System) to verify adequate tool access for eight fasteners which will have to be removed for OGS Water ORU Filter R&R during 17A. [After the recent Water ORU R&R successfully recovered the OGS from the delta pressure in the recirculation loop, the plan is to attempt a recovery maintenance task for the failed ORU. Mike found “no problem – all 8 of the 9/64 hex head fasteners on back of manifold can be accessed; need different tools for different fasteners.”]

FE-4 Bob Thirsk completed the weekly offloading of the WPA (Water Processor Assembly) from WRS (Water Recovery System) Rack 1 into a CWC-I (Collapsible Water Container-Iodine, #2004) with the common H2O transfer hose (which took about 23 min) from the PWD (Potable Water Dispenser) Auxiliary Port, then flushed the system.

Afterwards, Thirsk collected water samples from the ITCS LTL (Internal Thermal Control System Low Temperature Loop) and MTL (Moderate TL) in the Node-2 and from the MTL in the Lab, all for return to the ground on the Shuttle.

FE-5 Frank De Winne also collected water samples from the COL (Columbus Orbital Laboratory) ITCS for ground return.

Romanenko did the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

CDR Padalka completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM (Service Module). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The crew performed an abbreviated (1-hr) physical workout program on the CEVIS cycle ergometer (FE-3, FE-4), TVIS treadmill with vibration isolation (FE-1, FE-2, FE-5), and VELO cycle ergometer with bungee cord load trainer (CDR).

Afterwards, Frank transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

At ~6:15am EDT, the FE-5 had his weekly PFC (Private Family Conference), via S-band/audio and u-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop).

Sleeptime for the crew began at 10:00am, to last until 6:30pm tonight.

Weekly Science Update (Expedition Twenty -- Week 13)

3-D SPACE: No report.

AgCam (Agricultural Camera): No report.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): Standing by.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): No report.

BIOLAB (ESA): No report.

Biological Rhythms (JAXA, BIORHYTHMS): “Session #2 was completed on FD69. Data download was completed. We confirmed that the noise on a channel only continued 2 hours and did not impact science analysis. Session #2 was completed on FD71. Data download was completed.”

BISE (CSA, Bodies in the Space Environment): No report.

BISPHOSPHONATES: “Bob, thanks for completing your pill ingestion. Your next session is scheduled for 8/31.”

CARD (Long Term Microgravity: Model for Investigating Mechanisms of Heart Disease, ESA): No report.

CARDIOCOG-2: Complete.

CB (JAXA Clean Bench): No report.

CBEF (JAXA Cell Biology Experiment Facility): Plan temporary fan repair during 17A flight in order to start Space Seed no later than 9/6 or 9/10.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Tim, your last CCISS session is tied to the launch on FD2. We look forward to supporting your session.”

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack), MDCA/Flex: CIR resumed science operations this week. Methanol test point 2FLX192 was attempted however significant levels of gas bubbles in the fuel prevented a successful dispense and deployment. Heptane test point FLEX196 was also attempted. Gas bubbles again interfered with the dispense however the droplet was deployed and ignited although a significant velocity was impacted to droplet and the extinguish phenomena occurred outside the field of view of our imaging system CIR attributes the gas bubbles to dissolved CO2 from previous high pressure test points. Both the installed fuel reservoirs are now empty. A single heptane reservoir is in stowage and is planned to be installed. Additional fuel will arrive with the HTV along with fiber droplet tethers. A revised test matrix that orders test points from low pressure to high and the tethers are expected to improve future test success rates.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Completed.

Commercial 3 (JAXA): Completed.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

DOSIS (ESA): No report.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.

EDR (European Drawer Rack, ESA): The rack is continuously active in support of the Protein Crystallization Diagnostic Facility (PCDF) experiment. EDR is providing power/data and temperature control (via cooling loop) to PCDF.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

ENose (Electronic Nose): No report.

EPM (European Physiology Module): No report.

EPO (Educational Payload Operations, NASA): “Thank you Bob and Frank for completing a total of 8 EPO-Demos this week, which is a record for the most demos ever completed during a week. This past week’s efforts also assisted in setting an EPO record for the most demos completed during an increment totalling 14. You did not just give us quantity, you gave us quality. Fantastic job! We will make appropriate changes to future demos to make sure there is enough review and set-up time for the crew. Thank you for your time and dedication to educators and students worldwide.”

EPO 3-min Video (JAXA): No report.

EPO J-Astro Report (JAXA): Ongoing.

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): No report.

EPO Moon Score (JAXA): The JAXA Photo Moon team apologizes that the pre-calculated moon position was different from what you have actually seen from the ISS window. We are now investigating the Moon position simulation as well as downlinked photo.

EPO Try Zero-G (JAXA): “No report.

EPO Kibo Kids Tour (JAXA): Complete.

EPO Spiral Top (JAXA): No report.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): No report.

FACET (JAXA): No report.

FSL (Fluid Science Laboratory): No report.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter ECG (JAXA): No report.

HQPC (JAXA): To be launched by 34P.

ICE CRYSTAL (JAXA): Complete.

ICV (Integrated Cardiovascular): No report.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

INTEGRATED IMMUNE: No report.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): No report.

IRIS (Image Reversal in Space): No report.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): “Mike and Tim: Thanks for completing the final two LOCAD surface sampling sessions for this increment. Between the two of you, 15 sessions were performed! We have just a few swabbing kits remaining and we have been requested to look into possible water sampling so we want to reserve the remaining swabbing kits for that activity. You have collected a wealth of excellent and interesting data for us throughout this increment and today's results added to that. Thanks so much for the fantastic job!”

MEIS (Marangoni Experiment for ISS) in JAXA FPEF (Fluid Physics Experiment Facility): “Experiment is going very well. Hard drive (VRU) exchange by Crew member was conducted on GMT219 & 222. Up to now, we have successfully formed taller liquid bridge which length varied from 22.5 to 60 mm with 30 mm diameter. In each liquid bridge, the onset of oscillatory convection was able to be detected and the transition temperature difference to oscillation was determined precisely. W18 & W19: ; Experiment#3 (GMT 215) L/D=0.15 &0.18 ; Experiment#4(GMT 216) 0.20 &0.50 ; Experiment#5 (GMT 217) L/D=1.0; Experiment#6 (GMT 218) L/D=0.5; Experiment#7 (GMT 220) L/D=0.5; Experiment#8 (GMT 222) L/D=2.0; Experiment#9 (GMT 223) L/D=1,5; Experiment#10 (GMT 224) L/D=0.75; Experiment#11 (GMT225) L/D=2.0; Video Recoding Unit Hard Drive Exchange (GMT 226); Experiment#12 (GMT229) L/D=1.75; Experiment#13 (GMT230) L/D=1.25; Experiment#14 (GMT231) L/D=0.33; Experiment#15 (GMT231) L/D=0.63; Experiment#16 (GMT231) L/D=0.69 & 0.87; Video Recording Unit Hard Drive Exchange (GMT237); Completed the second series of Marangoni Experiment with Extra success. These data will contribute to the science of Oscillatory Marangoni Convection in Liquid Bridge. Appreciate all related efforts!”

MAXI (Monitor of All-sky X-ray Image, JAXA): Medium Rate Data Line checkout #3 for MAXI complete.(6 Aug); Attitude Determination System (ADS) C/O1 complete (8 Aug); All the 12 Gas Slit Cameras (GSC) have been turned on (13 Aug); Set Gas Slit Camera (GSC) protection (14 Aug); Upload alignment parameter. Start GSC commissioning (15 Aug); Cooling system of SSC on. CCD of GSC temp reached -60 degree (16 Aug); Initial set of SSC. GSC performing mostly as expected (17 Aug); Released the first light image of MAXI (18 Aug) http://maxi.riken.jp/top/; Continued several tunings.; On 25 Aug, all MAXI data transmission through Ethernet stopped. Troubleshooting is on-going. Plan to reboot MAXI system on 28 Aug.

MDCA/Flex: See under CIR.

Micro-G Clay (JAXA EPO): Complete.

MMA (JAXA/Microgravity Measurement Apparatus): MMA data downlink performed on 7/1. FSL Optical Target #4 check-out was completed by 7/1.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): No report.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): No report.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY: Ongoing.

PADLES (JAXA, Area PADLES 3; Passive Area Dosimeter for Lifescience Experiment in Space): Plan to retrieve on 17A. Removing for return is scheduled on FD2 of 17A.

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): No report.

PCG (JAXA, Protein Crystal Growth): “We started the experiment as scheduled. Crystallization is continuing in the PCG/PCRF automatically until PCG retrieval by 18S. To obtain the high quality protein crystals, 47 protein crystallizations will be conducted in microgravity environment for 74 days at 20 degrees C. Temperature monitoring by ground operation is now in progress. Experiment continues in good condition.”

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

POLCA/GRAVIGEN (ESA): Complete.

RadGene & LOH (JAXA): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (Space Environment Data Acquisition Equipment-Attached Payload, JAXA): ; Heavy Ion Telescope/ Plasma Monitor C/O complete (7 Aug); Neutron Monitor C/O Complete (8 Aug); Standard Dose Monitor C/O complete(12 Aug); Electronic Device Evaluation Equipment C/O complete (11 Aug); Electric Device Evaluation Equipment (EDEE) start (8/16); AOM: Atomic Oxygen Monitor function was verified (8/17); All data evaluation may take 2 more weeks.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Thanks for completing another week of sleep logging! We will be downlinking the data early next week. Mike, thanks for downloading all the Actiwatches last week. All the data looked good.”

SMILES (JAXA): Exposed Payload, to be launched by HTV1.

SOLAR (Solar Monitoring Observatory): No report.

SOLO (Sodium Loading in Microgravity): “Mike, we want to pass on to you very appreciative words of thanks for your participation in SOLO. All activities have been completed nominally for the SOLO diet session#2 and the science team looks very much forward to the analysis of the samples, logs and data. We hope you enjoy all the food and drinks of your choice now.”

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): “Mike and Tim, thank you for another very successful SPHERES session. Your dedication and commitment to SPHERES is greatly appreciated. We’re interested to hear your response to the questions regarding the Group A Test 5 that have been raised. We’re always interested in improving SPHERES wherever we can. You have provided much valuable science for us to analyze so we can continue to develop test sessions that push the limits!”

SPICE (Smoke Point In Co-flow Experiment): No report.

SWAB (Characterization of Microorganisms & Allergens in Spacecraft): “Mike, thanks for completing the first in-flight SWAB water collection operations! The expiration dates you noticed on the hardware had been reassessed on the ground some time ago and were deemed unneeded, so, we will update our words for future operations to address that issue. In regards to the sample mixing, the PI team is working on some ideas to ease that activity for the next ops and thank you very much for your description of the issue along with your explanation of how you made it work!”

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage). No report.

CEO (Crew Earth Observations): “We received your images of Hurricane Bill and thank you for your persistence in following this tropical system. Your recent views of the Black Point Lava Flow where the Lunar Electric Rover (LER) will be tested in the coming weeks have caused a great deal of excitement. Your detailed views of the area clearly show the road that the teams will travel on before they head out into the field. Team members will also be using your images for an overview briefing this week. Absolutely outstanding! Thank you for these images! Your striking image of Sevastopol, Ukraine was published on NASA/GSFC’s Earth Observatory website this past weekend. Your photograph highlighted the jagged coastline of the southern Crimean Peninsula and the various docking areas of Sevastopol. Thanks again for all the great images – we never get tired of looking at them!”

CEO photo targets uplinked for today were Hurricane Danny, off New England (predicted to reach hurricane strength by the time of this pass, looking left of track in early morning light for good definition of the upper surface), Sao Paulo-Rio de Janeiro Aerosol (aerosol opportunity as high pressure sets in over the Rio-Sao Paulo basin. Looking right down the coast and shooting high obliques towards the limb. Trying to include the coastline in some images), and Rio de Janeiro, Brazil (HMS Beagle site. Looking right for Rio’s prime visual cue, Guanabara Bay. Darwin undertook an expedition inland from Rio, presently a city of more than 7 million).

CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 8:46am EDT [= epoch])
Mean altitude -- 347.9 km
Apogee height – 354.0 km
Perigee height -- 341.8 km
Period -- 91.50 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0009105
Solar Beta Angle -- 39.7 deg (magnitude increasing)
Orbits per 24-hr. day -- 15.74
Mean altitude loss in the last 24 hours -- 44 m
Revolutions since FGB/Zarya launch (Nov. 98) -- 61750

STS-128/Discovery - 17A Crew & Mission Timeline:
v CDR: Fredrick W. “Rick” Sturckow
v PLT: Kevin A. Ford
v MSs: Patrick G. Forrester; Jose M. Hernandez; John D. “Danny” Olivas; Christer Fuglesang
v ISS FE-2s: Nicole Passonno Stott (UP); Timothy L. Kopra (DOWN)
· FD01 (8/28) -- Launch 11:59pm EDT
· FD02 (8/29) -- TPS inspection using OBSS; checkout EVA suits; prepare for rendezvous/docking
· FD03 (8/30) -- Rendezvous; RPM, docking (~9:03pm); exchange Soyuz seat liners
· FD04 (8/31) -- Unberth, install & ingress MPLM;; prepare for EVA1; campout (Olivas & Stott)
· FD05 (9/01) -- EVA1; remove ATA, retrieve EUTEF & MISSE 6; transfer racks
· FD06 (9/02) -- Focused TPS inspection with OBSS on SRMS; payload/cargo transfers; prepare for EVA2; campout (Olivas & Fuglesang)
· FD07 (9/03) -- EVA2; install new ATA; prep old ATA for return; cargo transfers
· FD08 (9/04) -- Off duty; prepare for EVA3; campout (Olivas & Fuglesang)
· FD09 (9/05) -- EVA3; deploy S3 PAS, R&R RGA 2, connect PMA-3 heater cables, R&R S0 RPCM, route Node-3 cabling
· FD10 (9/06) -- Crew off duty; final cargo transfers; initiate O2 transfer
· FD11 (9/07) -- Egress & uninstall MPLM; berth MPLM in PLB, terminate O2 transfer, close hatches
· FD12 (9/08) -- Undock from ISS (~3:27pm EDT); perform flyaround; dump H2O; late inspection
· FD13 (9/09) -- Cabin stow; Orbiter FCS checkout; RCS hot fire; H2O dump
· FD14 (9/10) -- Nominal deorbit (6:05pm); landing (~7:08pm KSC).

ISS Crew Sleep Shifting: To synchronize the ISS crew’s timeline with STS-128/17A arrival and docked activities, the station’s wake/sleep cycle is undergoing a number of shifts starting on 8/29. For the next few days, the schedule is as follows:

8/29
Sleep: 10:00a – 6:30p EDT
8/30
Sleep: 5:00a – 1:30p
8/31
Sleep: 5:00a – 1:30p
9/01
Sleep: 4:30a – 1:00p
9/02
Sleep: 4:00a – 12:30p
9/03
Sleep: 4:00a – 12:30p
9/04
Sleep: 3:30a – 12:00p
9/05
Sleep: 3:30a – 12:00p
9/06
Sleep: 3:00a – 11:30a
9/07
Sleep: 3:00a – 11:30a
9/08
Sleep: 2:30a – 11:00a
9/09
Sleep: 2:30a – 11:00a


Significant Events Ahead (all dates Eastern Time, some changes possible!):
08/30/09 -- STS-128/Discovery/17A docking – ~9:03pm EDT)
09/08/09 -- STS-128/Discovery/17A undocking – ~3:27pm
09/10/09 -- STS-128/Discovery/17A deorbit burn – ~6:05pm
09/10/09 -- STS-128/Discovery/17A landing (KSC; ~7:08pm)
09/10/09 -- H-IIB (JAXA HTV-1) launch (~1:04pm EDT)
09/16/09 -- H-IIB (JAXA HTV-1) berth w/SSRMS
09/29/09 -- Progress 34P undock
09/30/09 -- Soyuz TMA-16/20S launch
10/02/09 -- Soyuz TMA-16/20S docking (SM aft, until MRM-2 w/new port)
10/11/09 -- Soyuz TMA-14/18S undock
10/14/09 -- H-IIB (JAXA HTV-1) unberth
10/15/09 -- Progress 35P launch
11/10/09 -- 5R/MRM-2 (Russian Mini Research Module 2) on Soyuz-U
11/12/09 -- STS-129/Atlantis/ULF3 - ELC1, ELC2
12/07/09 -- Soyuz TMA-17/21S launch
12/26/09 -- Progress 36P launch
02/03/10 -- Progress 37P launch
02/04/10 -- STS-130/Endeavour/20A – Node-3 + Cupola
03/18/10 -- STS-131/Discovery/19A – MPLM(P), LMC
04/02/10 -- Soyuz TMA-18/22S launch
04/27/10 -- Progress 38P launch
05/14/10 -- STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1
05/29/10 -- Soyuz TMA-19/23S launch
06/25/10 -- Progress 39P launch
07/29/10 -- STS-133/Endeavour (ULF5 – ELC4, MPLM) or STS-134/Discovery (ULF6 – ELC3, AMS)
08/11/10 -- Progress 40P launch
09/16/10 -- STS-133/Endeavour (ULF5 – ELC4, MPLM) or STS-134/Discovery (ULF6 – ELC3, AMS)
09/29/10 -- Soyuz TMA-20/24S launch
10/19/10 -- Progress 41P launch
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton